Find the present values of the following cash flow streams at a6% discount rate. Do...

  1. Find the present values of the following cash flow streams at a 6% discount rate. Do not round intermediate calculations. Round your answers to the nearest cent.


    012345

    Stream A: $  
    Stream A$0$100$400$400$400$250
    Stream B$0$250$400$400$400$100

    Stream B: $  


  2. What are the PVs of the streams at a 0% discount rate? Round your answers to the nearest dollar.

    Stream A: $  

    Stream B: $  
Answer #1

Present value=Cash flows*Present value of discounting factor(rate%,time period)

At 6%:

A:

Present value=100/1.06+400/1.06^2+400/1.06^3+400/1.06^4+250/1.06^5

=$1289.84(APPROX).

B:

Present value=250/1.06+400/1.06^2+400/1.06^3+400/1.06^4+100/1.06^5

=$1319.26(APPROX).

At 0%:

A:

Present value=100+400+400+400+250

=$1550

B:

Present value=250+400+400+400+100

=$1550